what is the period of sec(1/4x)
Answer by lwsshak3(11628)

what is the period of sec(1/4x)
equation of sec function: y=sec(Bx)
period=2π/B
For given function, sec(1/4x):
B=1/4
period=2π/(1/4)=8π